OpenStudy (shamil98):

Graph the area bounded by y< 1/2x + 6 , x + 3y>= 12, x <= 0 , and x <=12.

ganeshie8 (ganeshie8):

familiar wid graphing lines ?

OpenStudy (shamil98):

Yeah, but not the bounded part ._.

ganeshie8 (ganeshie8):

good, i can help u wid bounding part :) first draw the corrsponding lines

ganeshie8 (ganeshie8):

y = 1/2x + 6 , x + 3y = 12, x = 0 , and x =12.

ganeshie8 (ganeshie8):

go thru this aswell :- http://www.purplemath.com/modules/syslneq.htm

OpenStudy (shamil98):

What does the x+3y >= 12 mean?

ganeshie8 (ganeshie8):

first graph the line, then we can **see** u wat does it mean :)

ganeshie8 (ganeshie8):

x+3y >= 12 corresponding line : x+3y = 12 can u graph this line ?

ganeshie8 (ganeshie8):



OpenStudy (shamil98):

Yes, it would be y = -1/3 x + 4

ganeshie8 (ganeshie8):

yup, can u draw it, you may "Reply use drawing" by hovering ur mouse pointer on top right of my drawing above

ganeshie8 (ganeshie8):